OpenStudy (anonymous):

Find the area enclosed by the x-axis and the curve x = 3 + e^t, y = t − t^2.

OpenStudy (anonymous):

I am not sure if it works, but it is worth trying From the equation \(x = 3 + e^t\) \[x = 3 + e^t\]\[x -3= e^t\]\[t \ln e = \ln (x-3)\]\[t= \ln (x-3)\] Sub t = ln (x-3) into the equation of y = t - t^2 \[y = \ln (x-3) + [\ln (x-3)]^2\] Sub y=0 into \(y = \ln (x-3) + [\ln (x-3)]^2\) to find the bounds: \[0 = \ln (x-3) + [\ln (x-3)]^2\]\(ln(x-3)= 0 \) or \(\ln (x-3) -1 =0\) Solve x from these two equations to find the bounds. Then, you can integrate the equation \(y = \ln (x-3) + [\ln (x-3)]^2\) to find the area.